# Fisher-Geometric Diffusion in Stochastic Gradient Descent: Optimal Rates, Oracle Complexity, and Information-Theoretic Limits

### 10 Apr 2026 v3

@@ 8.3 Claim 3: Scalar Temperature Matching Fails—Directionality Is the Signature of Geometry
Additional robustness experiments under misspecification and approximate exchangeability are reported in Appendix B.
@@ 9.1 Summary of Contributions
The paper’s central contribution is the identification of the mini-batch gradient covariance from sampling design and exchangeability as the projected gradient-noise matrix $G^{\star}(\theta)$—which reduces to Fisher information at $\theta^{\star}$ under correct specification and, for general losses, induces Godambe/sandwich geometry together with $H^{\star}$. This identification endogenizes the diffusion coefficient, determines the Lyapunov stationary covariance, and selects the natural metric for rates and complexity, yielding matching $1/N$ bounds governed by intrinsic dimension $d_{\mathrm{eff}}$ and condition number $\kappa_{F}$.
  • Proven: Alignment (Theorem 4.3), diffusion (Theorem 5.4), OU/Lyapunov (Corollary 5.7), discrete bridge (Proposition 5.9), matching bounds (Theorem 6.5 and Proposition 6.10), and oracle complexity (Theorem 7.3).
  • Sketched: Batch-size scaling (Corollary 7.8) and variance-reduction bridge (Remark 9.1).
  • Deferred: Misspecification, weak dependence, explicit batching policies, and OR-specific applications (Appendix B provides preliminary experiments).
9.2 Implications for Algorithm Design
In the diffusion regime the key knob is the effective temperature $\tau=\eta/b$. Changing $b$ does more than “reduce variance”; it changes the diffusion amplitude in a directional way determined by the identified geometry, and therefore changes the Lyapunov-predicted error floor. Because the intrinsic noise is matrix-valued, variance-reduction interventions should be evaluated by their effect on risk in a chosen metric (e.g., $\mathrm{Tr}(Q\Sigma)$) rather than Euclidean variance alone, connecting classical OR variance-reduction practice to SGD through the correct geometry.
@@ 9.2 Implications for Algorithm Design
SGD does not precondition the drift; natural-gradient methods do so by reweighting $\nabla L$. Our results explain why plain SGD nevertheless exhibits geometry-aware behavior through its noise structure: drift contraction and diffusion injection are distinct levers. The control view also suggests adaptive batching—increasing $b$ (cooling) when marginal variance reduction is valuable in the performance metric of interest, and decreasing it when exploration matters.
@@ 9.4 Estimating $G^{\star}$ in Practice
The preceding theory identifies $G^{\star}(\theta)$ as the central primitive. In applications where $G^{\star}$ is not available analytically, several estimation strategies provide operationally useful approximations. The simplest is the empirical covariance of per-sample gradients: given a pilot batch $\{z_{1},\ldots,z_{m}\}$ at $\theta_{t}$, the sample covariance $\hat{G}(\theta_{t}):=\frac{1}{m-1}\sum_{i=1}^{m}(\psi_{i}-\bar{\psi})(\psi_{i}-\bar{\psi})^{\top}$ (where $\psi_{i}:=\nabla_{\theta}\ell(\theta_{t};z_{i})$) is the natural plug-in estimator and converges at the standard $O(1/\sqrt{m})$ rate in operator norm. For large $d$, a diagonal approximation $\hat{G}_{\mathrm{diag}}$ retains the per-coordinate noise scales and suffices for estimating $d_{\mathrm{eff}}$ and detecting gross anisotropy; when the leading eigendirections of $G^{\star}$ are the primary target, a rank-$k$ approximation via streaming PCA on per-sample gradients requires only $O(dk)$ storage. If $G^{\star}(\theta)$ varies slowly along the optimization path, an exponential moving average $\hat{G}_{t}=(1-\alpha)\hat{G}_{t-1}+\alpha\,(\psi_{t}-\bar{\psi}_{t})(\psi_{t}-\bar{\psi}_{t})^{\top}$ tracks the local geometry with minimal overhead.
[…]
@@ 9.5 Limitations and Future Work
Appendix A Technical Lemmas and Concentration Tools
This appendix collects the auxiliary lemmas invoked by the main-text proofs. All spine results—the alignment theorem, the diffusion limit, the OU linearization, the Lyapunov corollary, the upper and lower bounds, and the oracle complexity guarantee—are proved in the body of the paper.
A.1 Finite-population covariance identity
The following identity, classical in survey sampling (Cochran, 1977; Serfling, 1980), is used in the proof of the alignment theorem (Theorem 4.3).
Lemma A.1 (Without-replacement covariance of a mini-batch mean).
Let $x_{1},\dots,x_{n}\in\mathbb{R}^{d}$ be deterministic vectors with mean $\bar{x}_{n}:=\frac{1}{n}\sum_{i=1}^{n}x_{i}$ and centered sample covariance
$S_{n}:=\frac{1}{n-1}\sum_{i=1}^{n}(x_{i}-\bar{x}_{n})(x_{i}-\bar{x}_{n})^{\top}.$
Let $B$ be a subset of $\{1,\dots,n\}$ of size $b$ drawn uniformly without replacement, and define $\bar{x}_{B}:=\frac{1}{b}\sum_{i\in B}x_{i}$. Then, conditional on $(x_{1},\dots,x_{n})$,
$\mathrm{Cov}(\bar{x}_{B}\mid x_{1:n})=\frac{1}{b}\left(1-\frac{b}{n}\right)S_{n}.$ (50)
Proof.
Write $\bar{x}_{B}=\frac{1}{b}\sum_{i=1}^{n}I_{i}x_{i}$ where $I_{i}=\mathbf{1}\{i\in B\}$. Under SRSWOR, $\mathbb{E}[I_{i}]=b/n$, $\mathrm{Var}(I_{i})=\frac{b}{n}(1-\frac{b}{n})$, and $\mathrm{Cov}(I_{i},I_{j})=-\frac{b(n-b)}{n^{2}(n-1)}$ for $i\neq j$. Expanding $\mathrm{Cov}(\bar{x}_{B}\mid x_{1:n})=b^{-2}\sum_{i,j}\mathrm{Cov}(I_{i},I_{j})x_{i}x_{j}^{\top}$ and collecting terms:
$b^{-2}\!\left[\frac{b(n-b)}{n^{2}}\sum_{i}x_{i}x_{i}^{\top}-\frac{b(n-b)}{n^{2}(n-1)}\sum_{i\neq j}x_{i}x_{j}^{\top}\right]$
$\quad=\frac{n-b}{bn^{2}}\cdot\frac{n}{n-1}\sum_{i=1}^{n}(x_{i}-\bar{x}_{n})(x_{i}-\bar{x}_{n})^{\top}=\frac{1}{b}\!\left(1-\frac{b}{n}\right)S_{n},$
where the last equality uses $(n-1)S_{n}=\sum(x_{i}-\bar{x}_{n})(x_{i}-\bar{x}_{n})^{\top}$. ∎
A.2 Score identities and information equality
Lemma A.2 (Score identities and information equality).
Assume $\theta\mapsto p_{\theta}(z)$ is differentiable, $\int p_{\theta}(z)\,dz=1$, and differentiation under the integral sign is valid. Then:
  1. (i)
    $\mathbb{E}_{\theta}[s_{\theta}(Z)]=0$, where $s_{\theta}(z):=\nabla_{\theta}\log p_{\theta}(z)$ is the score.
  2. (ii)
    $F(\theta)=\mathbb{E}_{\theta}[s_{\theta}(Z)s_{\theta}(Z)^{\top}]=-\mathbb{E}_{\theta}[\nabla_{\theta}^{2}\log p_{\theta}(Z)]$ (information equality).
Proof.
Standard; see, e.g., Lehmann and Casella (1998, Ch. 7). ∎
A.3 Lyapunov equation solution
@@
A.4 Product bound for Robbins–Monro weights
@@
A.5 Freedman’s inequality
Lemma A.5 (Freedman’s inequality).
Let $(M_{t})_{t\geq 0}$ be a martingale with $M_{0}=0$ and increments $\Delta M_{t}:=M_{t}-M_{t-1}$ satisfying $|\Delta M_{t}|\leq R$ almost surely. Let $V_{T}:=\sum_{t=1}^{T}\mathbb{E}[(\Delta M_{t})^{2}\mid\mathcal{F}_{t-1}]$ be the predictable quadratic variation. Then for any $v>0$ and $\lambda>0$:
$\Pr(M_{T}\geq\lambda\text{ and }V_{T}\leq v)\leq\exp\left(-\frac{\lambda^{2}}{2v+2R\lambda/3}\right).$
Equivalently, with probability at least $1-\delta$, whenever $V_{T}\leq v$:
$M_{T}\leq\sqrt{2v\log(1/\delta)}+\frac{2R}{3}\log(1/\delta).$
Proof.
See Freedman (1975) or Tropp (2015, Theorem 1.6). ∎
Appendix B Robustness Beyond Pure Exchangeability
The experiments in this appendix are exploratory robustness diagnostics, not part of the formal theoretical validation in Section 8. They demonstrate that the covariance-identification theory degrades gracefully under realistic departures from i.i.d. likelihood assumptions. Experiment A considers model misspecification, where curvature and noise geometry diverge; Experiment B models mild violations of exchangeability via variance inflation.
B.1 Experiment A: Misspecification (Probit DGP, Logistic Fit)
We generate n=50,000n=50{,}000 observations from a probit DGP and fit a logistic model, so that the optimization target is the KL projection $\beta^{\star}$ rather than the true $\beta_{0}$. Under misspecification the score covariance and curvature no longer coincide: $J^{\star}:=\mathrm{Var}\!\bigl(\nabla_{\beta}\ell(\beta;X,Y)\big|_{\beta=\beta^{\star}}\bigr)\neq H^{\star}:=\nabla^{2}\mathbb{E}[-\ell(\beta;X,Y)]\big|_{\beta=\beta^{\star}}$. The covariance-identification theorem (Theorem 4.3) nevertheless predicts $\mathrm{Cov}(g_{B}(\beta^{\star}))\approx(1/b)\,J^{\star}$, where $J^{\star}$ plays the role of $G^{\star}(\theta^{\star})$ in the general-loss notation. We estimate $J^{\star}$ from the full-sample scores and compare its diagonal entries to those of the empirical mini-batch covariance (batch size $b=256$, 800 replicates). Figure 9 shows close agreement (relative Frobenius difference $\approx 0.05$), confirming that the identification remains accurate under misspecification: mini-batch fluctuations are governed by the score covariance $J^{\star}$ rather than by curvature alone.
Figure 9: Experiment A: misspecification (probit DGP, logistic fit). Diagonal entries of the empirical mini-batch covariance vs. $(1/b)\,\hat{J}^{\star}$ (projected score covariance at $\beta^{\star}$); relative Frobenius difference $\approx 0.05$.
B.2 Experiment B: Approximate Exchangeability and Variance Inflation
@@
Appendix C Reproducibility and Code Availability
All experiments are reproducible from the accompanying code archive. The Python suite (Python_Codes/) includes a master script run_all.py that executes all experiments sequentially. Every script uses explicit random seeds (global seed =7=7) for deterministic reproducibility; small platform-level floating-point differences may produce minor variations ($\lesssim 1\%$) in Monte Carlo estimates.
Software environment.
Python $\geq$3.9; NumPy, SciPy, Matplotlib, Pandas. Tested under Python 3.10.12, NumPy 2.2, SciPy 1.15, Matplotlib 3.10. Total wall-clock time: $\approx$4 minutes on a single CPU core.
Script inventory.
Table 6 maps each script to the figures and tables it generates.
Table 6: Code inventory: scripts, outputs, and paper cross-references.
Script Experiment Figures / Tables Sec.
partA_batch_scaling.py 1 (Batch scaling) Fig. 2 8.1
partB_anisotropic_alignment.py 2 (Angle sweep) Fig. 3 8.1
partC_rotation_sweep.py 3 (Rotation) Figs. 57 8.3
bounds.py 4–6 (OU plateau, rate, directional) Tab. 4, Figs. 48, Tab. 5 8.18.3
partD_misspecification.py A (Misspecification) Fig. 9 App. B
partE_eps_exchangeable.py B ($\varepsilon$-exchangeable) Fig. 10 App. B
The script run_all.py executes all of the above in sequence. The script partF_optimal_batching.py produces additional exploratory figures not included in the paper.
Shared parameters ($d=10$ experiments).
Hessian eigenvalues log-spaced in $[1,100]$; $F^{\star}=H^{\star}$ (quadratic Gaussian model). Anisotropic gradient covariance $G^{\star}$ drawn as a random SPD matrix with eigenvalues in $[0.5,5]$ (seed =7=7); isotropic baseline $G_{\mathrm{iso}}=\sigma^{2}I$ with $\sigma^{2}=\mathrm{Tr}(G^{\star})/d$ (trace-matched). Constant-step experiments use $\eta=0.5/\lambda_{\max}(H)$, $b=64$, $\tau=\eta/b$; decaying-step experiments use $\eta_{t}=1.2/(t+50)$, $b=64$.
Verification.
All scripts were executed to completion with zero errors. Numerical outputs confirm: (i) log-log slopes $\approx-1$ for the $1/N$ risk decay; (ii) OU stationary variances match Lyapunov predictions to $1\%$; (iii) directional variance ratios align with the eigenstructure of $G^{\star}$; (iv) misspecification experiments show graceful degradation to the sandwich/Godambe prediction.

### 15 Apr 2026 v4

@@ 8.6 Claim 6: Fluctuation-Scale $(\eta/b)$-Collapse on the Raw Iterate Path
Additional robustness experiments under approximate exchangeability are reported in Appendix B.
@@ 9.1 Summary of Contributions
The paper’s central contribution is operational: we identify the mini-batch gradient covariance from the sampling design itself as $b^{-1}G^{\star}(\theta)$, converting the diffusion analysis of constant-step SGD from a modeling exercise into a statement about the replication budget. In likelihood models the identified matrix reduces to projected Fisher information; in general losses it pairs with $H^{\star}$ through the classical sandwich/Godambe geometry of M-estimation. At fixed batch size, the raw iterate path has a deterministic fluid limit (the gradient-flow ODE); the $\sqrt{b/\eta}$-scaled fluctuations converge to a linear diffusion with coefficient $G^{\star}$; near a nondegenerate optimum this is Ornstein–Uhlenbeck, and the corresponding Lyapunov covariance scaled by $\eta/b$ matches the stationary covariance of the linearized discrete recursion at leading order. In the local Fisher case, an upper bound of order $1/N$ is rate-order matched by an i.i.d. van Trees lower bound, with the intrinsic dimension $d_{\mathrm{eff}}$ and statistical condition number $\kappa_{F}$ appearing in the upper bound only.
9.2 Operational Implications for Sampling-Budget Design
The effective temperature $\tau=\eta/b$ is the single scalar that converts replication effort into diffusion amplitude, and the identification shows it does so in a directional way fixed by $G^{\star}$ rather than uniformly. For an OR practitioner, three consequences follow at the sampling-budget level. First, at a fixed oracle budget $N=Tb$, the batch size $b$ trades iterate count against per-step noise scale: increasing $b$ cools the error floor predicted by the linearized discrete Lyapunov benchmark (Proposition 5.13) in magnitude, but leaves the noise ellipsoid’s shape unchanged. Second, classical OR variance-reduction interventions (common variates, control variates, stratification) should be evaluated by their Loewner effect on $G^{\star}$, and thus by the induced reduction in $\mathrm{Tr}(Q\Sigma)$ for the operational metric $Q$, rather than by scalar component-variance alone—this is the natural bridge from classical variance-reduction practice to mini-batch SGD. Third, the directional amplification in Experiment 5 identifies exactly which parameter components absorb residual risk, and thus where additional sampling effort has the largest marginal return.
@@ 9.2 Operational Implications for Sampling-Budget Design
Plain SGD does not precondition the drift; natural-gradient methods do so by explicitly reweighting $\nabla L$. We do not claim plain SGD is “natural-gradient-like” in the drift sense; rather, its linearized local equilibrium benchmark (Proposition 5.13) reflects the noise covariance $G^{\star}$ through the Lyapunov balance. The identification also suggests adaptive batching—increasing $b$ (cooling) when marginal noise reduction is valuable, decreasing it when exploration matters.
@@ 9.4 Estimating $G^{\star}$ in Practice
Where $G^{\star}$ is not analytically available, the empirical covariance of per-sample gradients $\hat{G}(\theta_{t})=\frac{1}{m-1}\sum_{i}(\psi_{i}-\bar{\psi})(\psi_{i}-\bar{\psi})^{\top}$ is the natural plug-in and converges at the standard $O(1/\sqrt{m})$ rate in operator norm; for large $d$, a diagonal or rank-$k$ streaming approximation suffices for the quantities ($d_{\mathrm{eff}}$, dominant eigendirections, scale anisotropy) that enter the bounds. Given $\hat{G}$, one can diagnose whether isotropic temperature-matching is likely to fail (by checking whether $\hat{G}$ is far from a scalar multiple of $I$) and compare candidate batch sizes through the predicted stationary covariance $\Sigma_{\eta}$ of Proposition 5.13. A natural design heuristic (not proved here) is to regulate $\tau_{t}\propto 1/\lambda_{\max}(H(\theta_{t}))$ to avoid over-heating in stiff directions; when $d_{\operatorname{eff}}\ll d$, the linearized local equilibrium of Proposition 5.13 concentrates on the identifiable subspace captured by $\Pi_{\theta}$, which is where the identified geometry is nondegenerate.
@@ 9.5 Limitations and Future Work
The empirical design intentionally progresses from interpretable geometric diagnostics (direct covariance identification, 22D Lyapunov visualization) to quantitative raw-scale validation under direct SGD recursions in $d=10$, culminating in the fluctuation-scale $(\eta/b)$-collapse (Experiment 6; signature of Theorem 5.7) on the original iterate path.
Code availability.
All experiments in Section 8 and Appendix B are reproducible from the accompanying code archive (Python_Codes/), with master script run_all.py, explicit random seeds, and wall-clock under 5 minutes on a single CPU.
Appendix A Technical Lemmas
This appendix collects the auxiliary lemmas invoked by the main-text proofs. The finite-population covariance identity (Lemma 4.1) and score identities (Lemma 3.2) are stated and proved in the main text; we only restate here the auxiliary results (Lyapunov representation, Robbins–Monro product bound) that are invoked in later proofs but are not part of the main-text spine.
A.1 Lyapunov equation solution
@@
A.2 Product bound for Robbins–Monro weights
@@
Appendix B Robustness: A Stylized Variance-Inflation Perturbation
The following experiment probes one stylized perturbation of the fresh-sampling assumption of Theorem 4.3: a scalar variance-inflation factor $(1+\varepsilon)$ on the mini-batch covariance, modeling mild departures from ideal exchangeability (weak dependence, partial sample reuse). This is a specific perturbation diagnostic, not a general robustness theorem. A direct check of the identification itself under model misspecification is in Experiment 1 of Section 8.1.1.